OpenStudy (anonymous):

Let f(X)= 2x^2+x+1 , with x>=-1/4. Find the slope of the tangent line to the graph of y= f^(-1) (X) { f inverse of x } @ x =4

OpenStudy (anonymous):

do i first set my equation >= to -1/4 and solve it

ganeshie8 (ganeshie8):

you wanto find the slope at x = 4 for f^-1(x)

OpenStudy (anonymous):

ya

ganeshie8 (ganeshie8):

since 4 is x-coordinate of f^-1(x), 4 will be y-coordinate of f(x). lets find teh corresponding x-coordinate of f(x) :- 4 = 2x^2 + x + 1

ganeshie8 (ganeshie8):

By inspection x = 1 thus, (1, 4) is a point on f(x), which forces (4, 1) to be a point on f^-1(x)

ganeshie8 (ganeshie8):

next find \(f'(1)\) :- \(f(x) = 2x^2+x+1\) \(f'(x) = 4x+1\) \(f'(1) = 4(1) + 1 = 5\)

ganeshie8 (ganeshie8):

\(\large f^{-1}(4) = \frac{1}{f'(1)} = \frac{1}{5} = 0.2\) see if that makes more or less sense

ganeshie8 (ganeshie8):

for more explanation have a look at :- http://oregonstate.edu/instruct/mth251/cq/Stage6/Lesson/inverseDeriv.html

OpenStudy (anonymous):

but where did we use -1/4

ganeshie8 (ganeshie8):

good question :)

ganeshie8 (ganeshie8):

you need to look at graph, to *understand* why the constrain x >= -1/4 is given. let me show u quick

ganeshie8 (ganeshie8):